### v1

@@ 3.3.1 ADF method
where denotes an average, is the polarization angle difference between individual pairs of polarization vectors at position and , where and are the corresponding polarization angles at position and , respectively, and is number of pairs of vectors with a displacement of . At scales much smaller than the scale for a variation of the large-scale B-fields structure (typically of 5 seen by Planck observations), the total angular dispersion function can be expressed as:
@@
When , i.e., rad, one can obtain .
For each subregion (a) and (e) in Oph-A, we calculate the angular structure function using HAWC data at 154 m via Eq. 6 and then perform the fit to the quadratic function described in Eq. 7 using the Python curve-fit routine. We select data points that 13.6″50″ to avoid fitting to structures higher than the beam size and to better fit to smaller-scale structures. Top panels in Figure 7 show structure functions of polarization angles as a function of the distance of displacement toward subregions (a) and (e). From the fitted value of the fits, we derive values of and toward subregions (a) and (e), respectively. We find these values in good agreement with mean polarization angle dispersion values estimated using our approach described in Section 3.2.1, which are 9.2° and 14.3° in subregions (a) and (e), respectively. Toward subregions (a) and (e), mean gas volume densities are and cm-3 respectively, and their mean non-thermal FWHMs are 1.74 and 1.45 km s-1, respectively. We substitute new values together with the above mean values of and into Eq. 2 to re-calculate strengths in subregions (a) and (e), which are mG and mG, respectively. These values are in good agreement with strengths obtained using the DCF method, where values are in ranges of 0.32–1.55 mG and 0.15–0.97 mG toward subregions (a) and (e), respectively, with corresponding mean values of 0.84 and 0.37 mG.
@@ 3.3.2 ST method
The ST method is applicable for the compressible B-fields mode and thus is expected to provide a more accurate estimate of B-field strengths (Skalidis et al., 2021). Another advantage of the ST method is that this method accounts for the correlation between the fluctuation and mean components of the total B-field energy not to be zeros. Here, we measure the maps of strength () in band C and D using the formula as appeared in Skalidis & Tassis (2021):
We show the maps of across Oph-A in band C and D in Figure 16. We find that value ranges from mG at 89 m and mG at 154 m, appeared to be lower than . A median value of mG is seen in both bands. For a direct comparison with those using other methods, we estimate the mean value of at 154 m toward two subregions (a) and (e), which are 0.19 and 0.07 mG, respectively. These values are about 4.5 times lower than those estimated using the DCF method.
@@ 3.3.3 DMA method
We also re-calculate trengths in two subregions (a) and (e) using the DMA method, , using Eq. 10 following Lazarian et al. (2022):
@@
We calculate the structure function of the central velocity of HCO+(43) line toward subregions (a) and (e), using the similar form in Eq. 6 but for the central velocity , shown in middle panels in Figure 7. We then measure strengths in two subregions (a) and (e) via Eq. 10, using the information from the together with and the mean values of in the subregions. The as a function of displacement is shown in bottom panels of Figure 7 for subregions (a) and (e). It can be seen that the resulting in the range is likely flatten within their uncertainty. Thus, in each subregion we take the mean of in this range to be the average value, which are 0.34 mG and 0.12 mG in subregions (a) and (e), respectively. These values are slightly higher than the resulting estimated using the ST method, but at the low-end of that measured using the DCF method, i.e., 2–3 times less than the mean values estimated using the DCF method presented in Section 3.2. These differences are indeed expected since the DCF method somewhat overestimates the strength in the cloud. We will discuss this issue in Section 4.1.
@@ 3.4 Mass-to-flux ratio
The relative importance of B-fields compared to gravity in the cloud can be quantified by the mass-to-magnetic-flux ratio (or so-called mass-to-flux ratio), , which can be estimated in units of the critical ratio, (Nakano & Nakamura, 1978): . We measure in the entire region of Oph-A, using the formula introduced by Crutcher (2004):
@@
Due to the lack of measurements of B-field strengths along the line-of-sight () in the Oph-A cloud, we adopt the component as the total B-field strengths : . For the following analysis, we will employ only the map estimated using the DCF method and inferred from the polarimetric data in band D since it covers the larger area of the higher-density region in Oph-A than that from the data in band C.
The top panel in Figure 8 shows the spatial distribution of the mass-to-flux ratio in Oph-A. It reveals an increasing gradient of from the central and eastern parts of the clump to the outer regions, spanning from 0.03–2.56, with a median value of 0.26. The region exhibiting the lowest values of 1 is close to the OphS1 star, suggesting that the region is firmly magnetically sub-critical, i.e., B-fields are strong enough to prevent the gravitational collapse. This can also be explained by the influence of the strong winds and radiation fields from the OphS1 star which tend to push the material away to to its vicinity at the west side. The mass-to-flux ratio increases toward the outskirt regions of the Oph-S1 star, especially toward regions along the ridge of the dense cloud (i.e. center, north, and south-east parts of the map) where the gas column density is cm-2. These regions are magnetically super-critical and so, the self-gravitational collapse might have occurred there, resulting in several protostars and a single protostellar core identified in the region (Enoch et al., 2009; Evans et al., 2009; Connelley & Greene, 2010).
@@ 3.5 Alfvénic Mach number
where is given by Eq. 3, and is the Alfvénic speed driven by the B-fields (), both in units of km s-1. The factor in Eq. 12 assumes that turbulence is isotropic, and allows the calculation of the three-dimensional (3D) Alfvénic Mach numbers. We note that adding this factor will provide upper limits to the estimated .
The middle panel in Figure 8 shows the map of toward Oph-A using results of polarimetric data at 154 m and HCO+(4–3) as the gas tracer. The values are in the range from 0.1–1.5 across the cloud, with a median value of 0.38. Most of Oph-A shows ¡1 (magnetically sub-Alfvénic), suggesting the B-field dominates turbulence in the cloud. However, two small regions at the north and southwest of the map exhibit 1, indicating strong Alfvénic motions of the gas (magnetically supper-Alfvénic). These regions are also associated with the positions of a few prestellar and protostellar cores, and YSO candidates (Enoch et al., 2009; Evans et al., 2009; Connelley & Greene, 2010).
@@ 3.6 Plasma parameter
To compare the magnetic presure () and thermal pressure () in Oph-A, we construct the map of the plasma beta parameter (), given as:
where is the thermal sound speed measured in units of km s-1, assuming = (see Section 3.2.2).
The bottom panel in Figure 8 shows the map of the parameter toward Oph-A in band D. We obtain ¡1 across the map with a median value of 0.02, implying that the cloud is supported by strong magnetic pressure which is much more prominent than the one induced by the thermal motions of the gas.
@@ 3.7 Virial mass analysis
To probe the dynamical state of the Oph-A cloud, we calculate the virial parameter (), which is the ratio of the virial mass () and the isothermal mass () of the cloud. Here, we assume the cloud as a uniform density sphere with an effective radius of as a whole, having a mean mass density of . The isothermal mass, , considering the self-gravity of the cloud can be estimated as:
@@
where , , and are respectively the mean values of the Alfvénic velocity, thermal sound speed, and of non-thermal velocity dispersion toward the cloud, is the gravitational constant.
We calculate by selecting an ellipse covering the area where is measured. This ellipse is centered at (RA, Dec) = () with the semi-major and semi-minor axes are =120 and =75, respectively. The is then determined as (0.063 pc assuming the distance to Oph-A of 137 pc). Within this area, the mean gas volume density is cm-3, corresponding to the mean mass density . The corresponding mean values of velocities are: , , and . We use these mean values of the velocities, non-thermal velocity dispersion, and the gas volume density to calculate and in Eqs. 14 and 15, yielding 15. This shows that the cloud is gravitationally unbound and that B-fields and turbulence can support the cloud against self-gravity. This high value of obtained in the cloud overalls could explain the low efficiency of star formation in rho Oph-A though in a few regions the density is very high up to 105 or 106 cm-3. This is consistent with the lack of prestellar cores and protostellar sources in the central region, which is seen in the lower-density parts of the cloud only (see e.g., Figure 8). In order to study smaller scales across the clump, polarimetric observations with a higher spatial resolution would be required.
@@ 4.1 Morphology and strengths of magnetic field in Oph-A
The B-field orientation in Oph-A is well-ordered and predominantly perpendicular to the ridge of the dense cloud (Section 3.1). The feature is illustrated most clearly by the data at the longer wavelength (154 m). The data at 89 m better traces warmer dust in the lower-density region of Oph-A, which surrounds the Oph-S1 star.
The B-field properties measured using SOFIA/HAWC polarization data are qualitatively consistent with previous measurements of B-field orientations in Oph-A obtained at 850 m using SCUBA-2/POL–2 (Kwon et al., 2018) and SCUBA/SCUPOL (Matthews et al., 2009) on JCMT and at NIR wavelengths (JHK bands) using IRSF/SIRPOL (Kwon et al., 2015). In the POL–2 dataset, Oph-A was divided into 10 subregions, identified based on their B-field position angle and the polarization fraction, covering the entire extent of HAWC maps (see Figure 15). The highest density region is characterized by an average B-field position angle of 50°, which is also consistent with the NIR data toward the L1688 main cloud (Kwon et al., 2015). Table 1 shows the median B-field position angle () in each subregion which has been covered also by the FIR and sub-mm observations. The median B-field position angles from the 89, 154, and 850 m datasets are in rather good agreement, with the difference in values ranging from 3 to 19 (30%). These differences could be explained by the fact that the sub-mm data probe colder dust layers of the cloud, while the FIR data trace warmer layers. Indeed, such foreground layers were identified through the detection of the self-absorption features in the spectra of the [O i] line at 63 m (Liseau et al., 2006) and the [C ii] line at 158 m (Mookerjea et al., 2018).
@@
Kwon et al. (2018) calculated in three subregions associated with the densest parts of Oph-A using non-thermal velocity dispersion from the N2H+(1–0) line (André et al., 2007) and the sub-mm data at 850 m. In our study, we calculate pixel-by-pixel toward the entire Oph-A region. To facilitate comparisons with results obtained from the sub-mm data, we calculate the mean value of in two out of these three subregions, covered by the HAWC data (see Table 1). Mean values of using FIR data are lower than that obtained with the sub-mm data, which traces colder dust layers in the cloud characterized also by the higher densities. The HAWC data, on the other hand, essentially traces the warmer layers in the lower-density region. Also, Kwon et al. (2018) adopted values of the polarization angle dispersion as the standard deviation of the polarization angles, which are much smaller than those using the spread of the dust polarization angles (see Section 3.2.1). As a consequence, the obtained with sub-mm data is expected to be stronger than the one obtained with FIR data. We find that our estimate for Oph-A is typically higher than those in other molecular clouds, ranging from 50 to 400 G (for e.g., Coudé et al., 2019; Ngoc et al., 2021; Pattle et al., 2021; Ward-Thompson et al., 2023), but similar to those found in the DR21 (OH) region (Poidevin et al., 2013) or the OMC cloud in Orion A region (Pattle et al., 2017; Hwang et al., 2021) where of a few mG was measured.
@@
In this study, we adopt component as the total B-field strength. If we consider the contribution of the LOS component, the total B-field strength will be stronger than what we measured. The mass-to-flux ratio, Mach Alfvénic number, and plasma parameter will decrease, but this will not significantly change our conclusion that B-fields are generally dominant over gravity, turbulent, and gas kinetic energies, especially toward the center of the cloud. Furthermore, as presented in Section 3.3, we test that the strengths obtained with the DCF method in subregions (a) and (e) are rather in good agreement with values obtained with the ADF method but higher than those using the ST and DMA methods (see Table 2). We use new values estimated with ADF, ST, and DMA methods toward two subregions (a) and (e) to estimate the mass-to-flux ratio, Mach Alfvénic number, and plasma, which are all higher than that measured with the DCF method (see Table 2). Subregion (a) changes from magnetically sub/trans-critical to super-critical but remains to be super-Alfvénic, and has ¡1 in overall. A similar behavior is seen in subregion (e) except that this subregion changes from trans- to super-Alfvénic (see Table 2).
@@
We now assume that the B-field strength in the entire cloud decreases by a factor of 4.5 as measured with the ST method. In this case, the majority of the cloud will become magnetically trans/super-critical (1), except for the regions in the vicinity of the Oph–S1 star (see the top panel of Figure 20). However, this trend of remains similar as obtained using . The cloud will become mostly super-Afvénic with ¿1, except for the southeast region of the cloud (see the middle panel of Figure 20). This assumption will also result in the Alfvénic velocity decreasing by a factor of 4.5 and thus, . This implies that the cloud is in the magneto hydrostatic equilibrium and may undergo gravitational collapse due to turbulent fluctuations on smaller scales.

### v2

@@ 3.3.1 ADF method
where denotes an average, is the polarization angle difference between individual pairs of polarization vectors at position and , where and are the corresponding polarization angles at position and , respectively, and is number of pairs of vectors with a displacement of . At scales much smaller than the scale for a variation of the large-scale B-field structure (typically of 5 seen by Planck observations), the total angular dispersion function can be expressed as:
@@
When , i.e., rad, we can obtain .
For each subregion (a) and (e) in Oph-A, we calculate the angular structure function using HAWC data at 154 m via Eq. 6 and then perform the fit to the quadratic function described in Eq. 7 using the Python curve-fit routine. We select data points that 13.6″50″ to avoid fitting to structures higher than the beam size and to better fit to smaller-scale structures. Top panels in Fig. 7 show structure functions of polarization angles as a function of the distance of displacement toward subregions (a) and (e). From the fitted value of the fits, we derive values of and toward subregions (a) and (e), respectively. We find these values in good agreement with mean polarization angle dispersion values estimated using our approach described in Sect. 3.2.1, which are 9.2° and 14.3° in subregions (a) and (e), respectively. Toward subregions (a) and (e), mean gas volume densities are and cm-3 respectively, and their mean non-thermal FWHMs are 1.74 and 1.45 km s-1, respectively. We substitute new values together with the above mean values of and into Eq. 2 to calculate strengths in subregions (a) and (e), which are mG and mG, respectively. These values are in good agreement with strengths obtained using the DCF method, where values are in ranges of 0.32–1.55 mG and 0.15–0.97 mG toward subregions (a) and (e), respectively, with corresponding mean values of 0.84 and 0.37 mG.
@@ 3.3.2 ST method
The ST method is applicable for the compressible B-fields mode and thus is expected to provide a more accurate estimate of B-field strengths (Skalidis et al., 2021). Another advantage of the ST method is that this method accounts for the correlation between the fluctuation and mean components of the total B-field energy not to be zeros. Here, we measure maps of strength () in band C and band D using the formula as appeared in Skalidis & Tassis (2021):
We show the maps of across Oph-A in band C and band D in Fig. 17. We find that value ranges from mG at 89 m and mG at 154 m, appeared to be lower than . A median value of mG is seen in both bands. For a direct comparison with those using other methods, we estimate the mean value of at 154 m toward two subregions (a) and (e), which are 0.19 and 0.07 mG, respectively. These values are about 4.5 times lower than those estimated using the DCF method.
@@ 3.3.3 DMA method
We also calculate trengths in two subregions (a) and (e) using the DMA method, , using Eq. 10 following Lazarian et al. (2022):
@@
We calculate the structure function of the central velocity of HCO+(43) line toward subregions (a) and (e), using the similar form in Eq. 6 but for the central velocity , shown in middle panels in Fig. 7. We then measure strengths in two subregions (a) and (e) via Eq. 10, using the information from the together with and the mean values of in the subregions. The as a function of displacement is shown in bottom panels of Fig. 7 for subregions (a) and (e). It can be seen that the resulting in the range is likely flatten within their uncertainty. Thus, in each subregion we take the mean of in this range to be the average value, which are 0.34 mG and 0.12 mG in subregions (a) and (e), respectively. These values are slightly higher than the resulting estimated using the ST method, but they are at the low-end of that measured using the DCF method (i.e., 2–3 times less than the mean values estimated using the DCF method presented in Sect. 3.2). These differences are indeed expected since the DCF method somewhat overestimates the strength in the cloud. We discuss this issue further in Sect. 4.1.
@@ 3.4 Mass-to-flux ratio
The relative importance of B-fields compared to gravity in the cloud can be quantified by the mass-to-flux ratio, , which can be estimated in units of the critical ratio, (Nakano & Nakamura, 1978): . We measure in the entire region of Oph-A, using the formula introduced by Crutcher (2004):
@@
Due to the lack of component measurements in the Oph-A cloud, we adopted the component as the total B-field strengths : . For the analysis, we employed only the map estimated using the DCF method and inferred from the polarimetric data in band D; this is because it covers a larger area of the higher density region in Oph-A than that coming from the data in band C.
The top panel in Fig. 8 shows the spatial distribution of the mass-to-flux ratio in Oph-A. It reveals an increasing gradient of from the central and eastern parts of the clump to the outer regions, spanning from 0.03–2.56, with a median value of 0.26. The region exhibiting the lowest values of 1 is close to the OphS1 star, suggesting that the region is firmly magnetically sub-critical, namely, B-fields are strong enough to prevent the gravitational collapse. This can also be explained by the influence of the strong winds and radiation fields from the OphS1 star which tend to push the material away to to its vicinity at the west side. The mass-to-flux ratio increases toward the outskirt regions of the Oph-S1 star, especially toward regions along the ridge of the dense cloud (i.e. center, north, and southeast parts of the map) where the gas column density is cm-2. These regions are magnetically super-critical and so, the self-gravitational collapse might have occurred there, resulting in several protostars and a single protostellar core identified in the region (Enoch et al., 2009; Evans et al., 2009; Connelley & Greene, 2010).
@@ 3.5 Alfvénic Mach number
where is given by Eq. 3, and is the Alfvénic speed driven by the B-fields (), both in units of km s-1. The factor in Eq. 12 assumes that turbulence is isotropic, and allows the calculation of the 3D Alfvénic Mach numbers. We note that adding this factor will provide upper limits to the estimated .
The middle panel in Fig. 8 shows the map of toward Oph-A using results of polarimetric data at 154 m and HCO+(4–3) as the gas tracer. The values are in the range from 0.1–1.5 across the cloud, with a median value of 0.38. Most of Oph-A shows ¡1 (magnetically sub-Alfvénic), suggesting the B-field dominates turbulence in the cloud. However, two small regions at the north and southwest of the map exhibit 1, indicating strong Alfvénic motions of the gas (magnetically super-Alfvénic). These regions are also associated with the positions of a few prestellar and protostellar cores, and YSO candidates (Enoch et al., 2009; Evans et al., 2009; Connelley & Greene, 2010).
@@ 3.6 Plasma parameter
To compare the magnetic presure () and thermal pressure () in Oph-A, we constructed the map of the plasma beta parameter (), given as:
where is the thermal sound speed measured in units of km s-1, assuming = (see Sect. 3.2.2).
The bottom panel in Fig. 8 shows the map of the parameter toward Oph-A in band D. We obtained ¡1 across the map with a median value of 0.02. This result implies that the cloud is supported by strong magnetic pressure, which is much more prominent than that induced by the thermal motions of the gas.
@@ 3.7 Virial mass analysis
To probe the dynamical state of the Oph-A cloud, we calculated the virial parameter (), which is the ratio of the virial mass () and the isothermal mass () of the cloud. Here, we assume the cloud as a uniform density sphere with an effective radius of as a whole, having a mean mass density of . The isothermal mass, , considering the self-gravity of the cloud can be estimated as:
@@
where , , and are the mean values of the Alfvénic velocity, thermal sound speed, and of non-thermal velocity dispersion toward the cloud, respectively, while is the gravitational constant.
We calculate by selecting an ellipse covering the area where was measured. This ellipse is centered at (RA, Dec) = () with the semi-major and semi-minor axes are =120 and =75, respectively. The is then determined as (0.063 pc assuming the distance to Oph-A of 137 pc). Within this area, the mean gas volume density is cm-3, corresponding to the mean mass density . The corresponding mean values of velocities are: , , and . We used these mean values of the velocities, non-thermal velocity dispersion, and the gas volume density to calculate and in Eqs. 14 and 15, yielding 15. This shows that the cloud is gravitationally unbound and that B-fields and turbulence can support the cloud against self-gravity. This high value of obtained in the cloud overalls could explain the low efficiency of star formation in Oph-A; however, in a few regions, the density is very high (up to 105 or 106 cm-3). This is consistent with the lack of prestellar cores and protostellar sources in the central region, which can be seen in the lower density parts of the cloud only (see e.g., Fig. 8). To study smaller scales across the clump, polarimetric observations with a higher spatial resolution would be required.
@@ 4.1 Morphology and strengths of magnetic field in Oph-A
The B-field orientation in Oph-A is well-ordered and predominantly perpendicular to the ridge of the dense cloud (Sect. 3.1). The feature is illustrated most clearly by the data at the longer wavelength (154 m). The data at 89 m better traces warmer dust in the lower density region of Oph-A, which surrounds the Oph-S1 star.
The B-field properties measured using SOFIA/HAWC polarization data are qualitatively consistent with previous measurements of B-field orientations in Oph-A obtained at 850 m using SCUBA-2/POL-2 (Kwon et al., 2018) and SCUBA/SCUPOL (Matthews et al., 2009) on JCMT and at NIR wavelengths (JHK bands) using IRSF/SIRPOL (Kwon et al., 2015). In the POL-2 dataset, Oph-A was divided into ten subregions, identified based on their B-field position angle and the polarization fraction, covering the entire extent of HAWC maps (see Fig. 15). The highest density region is characterized by an average B-field position angle of 50°, which is also consistent with the NIR data toward the L1688 main cloud (Kwon et al., 2015). Table 1 shows the median B-field position angle () in each subregion, which has been covered by the FIR and sub-mm observations a well. The median B-field position angles from the 89, 154, and 850 m datasets are in rather good agreement, with the difference in values ranging from 3 to 19 (30%). These differences could be explained by the fact that the sub-mm data probe colder dust layers of the cloud, while the FIR data trace warmer layers. Indeed, such foreground layers were identified through the detection of the self-absorption features in the spectra of the [O i] line at 63 m (Liseau et al., 2006) and the [C ii] line at 158 m (Mookerjea et al., 2018).
@@
Kwon et al. (2018) calculated in three subregions associated with the densest parts of Oph-A using non-thermal velocity dispersion from the N2H+(1–0) line (André et al., 2007) and the sub-mm data at 850 m. In our study, we calculated the pixel-by-pixel toward the entire Oph-A region. To facilitate comparisons with results obtained from the sub-mm data, we calculated the mean value of in two out of these three subregions, covered by the HAWC data (see Table 1). Mean values of using FIR data are lower than that obtained with the sub-mm data, which traces colder dust layers in the cloud characterized also by the higher densities. The HAWC data, on the other hand, essentially traces the warmer layers in the lower density region. In addition, Kwon et al. (2018) adopted values of the polarization angle dispersion as the standard deviation of the polarization angles, which are much smaller than those using the spread of the dust polarization angles (see Sect. 3.2.1). As a consequence, the obtained with sub-mm data is expected to be stronger than the one obtained with FIR data. We find that our estimate for Oph-A is typically higher than those in other molecular clouds, ranging from 50 to 400 G (for e.g., Coudé et al., 2019; Ngoc et al., 2021; Pattle et al., 2021; Ward-Thompson et al., 2023), but similar to those found in the DR21 (OH) region (Poidevin et al., 2013) or the OMC cloud in Orion A region (Pattle et al., 2017; Hwang et al., 2021) where of a few mG was measured.
@@
In this study, we adopt component as the total B-field strength. If we consider the contribution of the LOS component, the total B-field strength will be stronger than what we measured. The mass-to-flux ratio, Mach Alfvénic number, and plasma parameter will decrease, but this will not significantly change our conclusion that B-fields are generally dominant over gravity, turbulent, and gas kinetic energies, especially toward the center of the cloud. Furthermore, as presented in Sect. 3.3, we test that the strengths obtained with the DCF method in subregions (a) and (e) are rather in good agreement with values obtained with the ADF method but higher than those using the ST and DMA methods (see Table 2). We use new values estimated with ADF, ST, and DMA methods toward two subregions (a) and (e) to estimate the mass-to-flux ratio, Mach Alfvénic number, and plasma, which are all higher than that measured with the DCF method (see Table 2). Subregion (a) changes from magnetically sub/trans-critical to super-critical but remains to be super-Alfvénic, and has ¡1 in overall. A similar behavior is seen in subregion (e) except that this subregion changes from trans- to super-Alfvénic (see Table 2).
@@
We go on to assume that the B-field strength in the entire cloud decreases by a factor of 4.5, as measured with the ST method. In this case, the majority of the cloud will become magnetically trans- or super-critical (1), except for the regions in the vicinity of the Oph–S1 star (see the top panel of Fig. 20). However, this trend of remains similar as obtained using . The cloud will become mostly super-Afvénic with ¿1, except for the southeast region of the cloud (see the middle panel of Fig. 20). This assumption will also result in the Alfvénic velocity decreasing by a factor of 4.5 and thus, . This implies that the cloud is in the magneto hydrostatic equilibrium and may undergo gravitational collapse due to turbulent fluctuations on smaller scales.